# Project Page: https://jbruner23.github.io/consistyle/

### v1

@@ Abstract.
In text-to-image models, consistent character generation is the task of achieving text alignment while maintaining the subject’s appearance across different prompts. However, since style and appearance are often entangled, the existing methods struggle to preserve consistent subject characteristics while adhering to varying style prompts. Current approaches for consistent text-to-image generation typically rely on large-scale fine-tuning on curated image sets or per-subject optimization, which either fail to generalize across prompts or do not align well with textual descriptions. Meanwhile, training-free methods often fail to maintain subject consistency across different styles. In this work, we introduce a training-free method that achieves both style alignment and subject consistency. The attention matrices are manipulated such that Queries and Keys are obtained from the anchor image(s) that are used to define the subject, while the Values are imported from a parallel copy that is not subject-anchored. Additionally, cross-image components are added to the self-attention mechanism by expanding the Key and Value matrices. To do without shifting from the target style, we align the statistics of the Value matrices. As is demonstrated in a comprehensive battery of qualitative and quantitative experiments, our method effectively decouples style from subject appearance and enables faithful generation of text-aligned images with consistent characters across diverse styles.
journal: TOG
@@ 1. Introduction
As far as we can ascertain, this is the first method to decouple style from identity while ensuring diversity and prompt alignment, without requiring any subject-specific training. Our empirical evaluations demonstrate that our method consistently outperforms prior approaches in both style and prompt alignment, while maintaining subject consistency comparable to existing methods.
@@ 3.1. Method Overview
(1) Initial generation   We first run a vanilla generation using the SDXL model. During this pass, we store the intermediate value features $V_{sd}$, which capture the fine-grained texture and color details needed for consistent style preservation in subsequent generations.
(2) Correspondence computation   Next, we run a generation to compute a set of DIFT features (Tang et al., 2023), which are used to establish a correspondence mapping $C_{\alpha}$ for the subject indices $s_{\alpha}$ in each image $\alpha$ and only includes attention crossing, a component which allows images to attend to subjects of other images in the self-attention layer and does not rely on the correspondence mapping. The subject indices are obtained from the cross-attention layer using a threshold over the attention map matching the subject token query. The images generated for computing DIFT do not employ the $Q$ and $K$ modifications of Sec. 3.2 since these require the subject location information that is computed using DIFT. However, the component crossing of Sec. 3.3, in which $K$ and $V$ are enriched with key and value pairs from other images in the batch after applying AdaIn is applied to obtain some level of subject consistency while maintaining style.
(3) Final generation   In the final pass, we perform a full generation that integrates all components: i. Value-Preservation: We reuse the stored values from the initial vanilla run, maintaining stylistic consistency, ii. Attention Transfer: During the initial phase, we employ the correspondence mapping to inject query and key features, for aligning the subject details across images, and iii. Attention Crossing: Throughout the final generation, we apply attention crossing to allow image queries to attend to the Keys and Values of other images in the batch, which improves subject consistency, while using AdaIN to prevent style-leak between images due to the different Value distributions.
Figure 4. Qualitative comparison of our method with Consistory, DB-LoRA, and IP-Adapter demonstrates its effectiveness across varying text descriptions, character consistency, and style alignment. Unlike other methods, our approach preserves character features and maintains consistent appearance while faithfully adhering to the specified style and textual descriptions.
[…]
@@
Table 2. Comparing our method to other zero-shot methods that achieve style or identity consistency, each with its own distinct goal. The component modification part shows the modification of the presentation of each generated image, while the component import part shows how the self-attention of the model is modified to have a cross attention component. The methods in the table are Consistory (Tewel et al., 2024), Cross-Image Attention (Alaluf et al., 2023), StyleAligned (Hertz et al., 2023) and IlluSign (Bruner et al., 2025). Our method is the only one that has an identity goal as well as a style goal, on top of the prompt faithfulness goal, which requires a much more nuanced solution. Promppt-to-Prompt (Hertz et al., 2022) is not listed as its attention modifications are done in the cross-attention layer and not the self-attention layer.
@@ 3.2. Transferring Style While Maintaining Appearance
We aim to transfer style from the target image while adopting the semantic content of the source images. To achieve this, we first perform a vanilla pass through the SDXL model, during which we store the value matrices $V_{s}d$ from the self-attention layer of the decoder at the highest-resolution transformer block $[64\times 64]$, and only during the early diffusion steps, specifically at steps $\frac{n}{10}$ to $3\frac{n}{10}$ (where $n$ is the number of steps). We also obtain subject masks $s_{1},s_{2},\dots$ using a threshold over the attention maps matching the subject token query in the cross-attention layer.
@@ 3.3. AdaIN for Style Preservation in Attention Crossing
Our approach mitigates this by applying adaptive instance normalization (AdaIN) to the values before cross-subject attention, effectively isolating the semantic content from style-specific features. The $\mathcal{A}$ operator operation is defined as follows,
@@
where $K_{i},K_{j},V_{i},V_{j}$ are the keys and values in the self-attention layer matching $i,j$ images, and $s_{i},s_{j}$ are the mask indices of the subjects in images $i,j$ respectively.
@@ 3.4. Summarizing the differences from other methods
With a clear view of the method and the associated terminology, we revisit the comparison to previous work, now on a clear technical level. A comparison to the most similar contributions can be found in Table  2. The table separates the modification of the self-attention mechanism from the step of selectively importing content from other attention maps, creating a cross-attention scheme.
We observe notable differences in the handling of self-attention component imports and modifications across various methods. Most approaches rely on direct Value imports from the anchor image. For instance, StyleAligned and Consistory both utilize direct Value imports, with Consistory further incorporating hidden state injections for enhanced visual consistency. Cross-Image Attention (Alaluf et al., 2023) similarly relies on direct Value modification, emphasizing precise texture and color transfer.
In contrast, our approach avoids direct Value imports to prevent the style misalignment typically associated with direct appearance transfer. Instead, we employ targeted Value modifications, aligning $V_{sd}$ to the original $SDXL$ features for improved style alignment. Additionally, we apply AdaIN to regulate the statistical properties of imported Values, ensuring smoother integration into the target domain. Unlike Consistory and Cross-Image Attention, we do not use the hidden values of the feature embedding for the self-attention intervention, and instead modify $Q$, $K$, and $V$ selectively, giving $V$ a different treatment. StyleAligned modified only $Q$ and $K$, using AdaIn that in our method is applied for the $V$ of the dictionary expansion part (the crossing of attentions).
There are also several differences, which are not captured in Table  2. For example, the timing of the attention intervention varies across methods. Key and Query modifications as well as the $V_{sd}$ injection in Consistyle, are confined to early generation steps, while other modifications and imports are applied across all time steps as specified in Sec. 3.2. Also, the Keys and Queries modification is done only on the text-guided part of the batch in the self-attention layer, where the $V_{sd}$ injection and attention crossing components are also applied on the non-guided elements, which seems to yield a slight improvement.
@@ 4. Experiments
We evaluate our method using three core metrics: prompt alignment, consistency, and style alignment, comparing it against three state-of-the-art (SOTA) baselines designed for consistent image generation. The first baseline is a training-free Consistory approach (Tewel et al., 2024), the second is an encoder-based method, IP-Adapter (Ye et al., 2023), and the third is a personalization training-based method, DreamBooth-LoRA (DB-LoRA) (Ryu, 2023). Unlike Consistory, which directly modifies the diffusion process, both IP-Adapter and DB-LoRA utilize a single reference image for personalization.
@@ 4.1. Qualitative Results
Our method is designed to accommodate a wide range of visual styles, including highly detailed photographic renderings, abstract illustrations and 3D aesthetics. In Fig.  4 and Fig.  ConsiStyle: Style Diversity in Training-Free Consistent T2I Generation we illustrate its ability to preserve style, maintain character consistency, and align with textual descriptions. As shown, IP-Adapter and DB-LoRA exhibit strong character consistency, yet they fail to respect the intended style and often feature very little variation. Notably, in both subjects, IP-Adapter and DB-LoRA render the kitten using a realistic style and the dragon with a 3D animation style across all images, disregarding the specified stylistic variations due to overfitting to the reference image. Consistory demonstrates strong character consistency and faithful prompt alignment; however, it frequently fails to harmonize the subject with the stylistic setting. In Fig.  4, this issue is evident in the dragon images for the film noir and line-art styles—both intended to be black and white, yet Consistory generates colored images or subjects. This misalignment extends beyond color. For example, in the kitten images (first and fourth row), the subject exhibits a highly realistic texture that clashes with the surrounding stylistic context. Furthermore, as shown in Fig.  5, although the compositions may initially seem coherent, closer examination reveals integration flaws—characters often appear visually ‘stitched’ onto the background rather than naturally embedded within the scene. In contrast, our approach maintains both consistency and stylistic harmony, allowing characters to seamlessly integrate into their environments and preserving the intended aesthetic.
The application of our method to multiple anchor images is demonstrated in Fig. 7.
@@ 4.2. Quantitative Results
Table 3. Numerical Results compared to the baselines. (Mean ± SD)
Method DreamSim $\downarrow$ CLIPScore $\uparrow$ CLIPScore, Styled $\uparrow$ LPIPS $\downarrow$ Gram L2 $\downarrow$
Consistyle (ours) $0.40\pm 0.10$ $\mathbf{32.84\pm 1.66}$ $\mathbf{36.03\pm 1.69}$ $\mathbf{0.21\pm 0.06}$ $\mathbf{1.25\pm 0.69}$
Consistory $0.33\pm 0.12$ $32.75\pm 1.53$ $35.34\pm 1.67$ $0.27\pm 0.07$ $1.85\pm 1.13$
IP Adapter $\mathbf{0.25\pm 0.08}$ $30.98\pm 2.03$ $32.07\pm 2.08$ $0.43\pm 0.07$ $2.96\pm 1.57$
DreamBooth-LoRA $0.28\pm 0.13$ $32.43\pm 1.74$ $34.33\pm 2.20$ $0.42\pm 0.07$ $2.6\pm 1.32$
To evaluate our approach we generated a comprehensive dataset of 25 prompt groups, each containing a subject description paired with ten distinct prompts. Additionally, we curated two style groups, each containing ten diverse styles spanning categories such as black-and-white, 3D, illustration, and photographic. Combining each prompt group with each style group resulted in 50 unique sets, each being a separate experiment consisting of a batch of $B=10$ images, which differ in both prompt and style (no prompt and no style are repeated twice).
For our evaluation we employ several automated metrics following prior work (Tewel et al., 2024; Junyao et al., 2024; Avrahami et al., 2024). Text alignment is measured using CLIPScore (Hessel et al., 2021) , including with and without style descriptions. Subject consistency is evaluated using DreamSim (Fu et al., 2023) as proposed by (Tewel et al., 2024) with background removal, and style alignment is assessed by LPIPS (Zhang et al., 2018) and Gram Matrix distance (Gatys et al., 2015) against vanilla SDXL images as style references.
The results of the experiments are presented in Table  3. As can be seen, our method outperforms all of the baselines in the prompt alignment and style alignment scores. However, we note a critical limitation in automated metrics for consistency when operating in style-diverse settings. Automated methods tend to over-penalize stylistic variation, potentially rewarding overly consistent outputs that lack stylistic diversity.

### v2

# ConsiStyle: Style Diversity in Training-Free Consistent T2I Generation

@@ Abstract.
In text-to-image models, consistent character generation is the task of achieving text alignment while maintaining the subject’s appearance across different prompts. However, since style and appearance are often entangled, the existing methods struggle to preserve consistent subject characteristics while adhering to varying style prompts. Current approaches for consistent text-to-image generation typically rely on large-scale fine-tuning on curated image sets or per-subject optimization, which either fail to generalize across prompts or do not align well with textual descriptions. Meanwhile, training-free methods often fail to maintain subject consistency across different styles. In this work, we introduce a training-free method that, for the first time, jointly achieves style preservation and subject consistency across varied styles. The attention matrices are manipulated such that Queries and Keys are obtained from the anchor image(s) that are used to define the subject, while the Values are imported from a parallel copy that is not subject-anchored. Additionally, cross-image components are added to the self-attention mechanism by expanding the Key and Value matrices. To do without shifting from the target style, we align the statistics of the Value matrices. As is demonstrated in a comprehensive battery of qualitative and quantitative experiments, our method effectively decouples style from subject appearance and enables faithful generation of text-aligned images with consistent characters across diverse styles.
Code will be available at our project page: jbruner23.github.io/consistyle.
journal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6article: 263publicationmonth: 12doi: 10.1145/3763303copyright: noneccs: Computing methodologies Computer graphicsccs: Computing methodologies Machine learning
This is the author version of a paper accepted to SIGGRAPH Asia 2025.
The final published version appears in ACM Transactions on Graphics.
@@ 1. Introduction
As far as we can ascertain, this is the first training-free method to jointly decouple style from identity while ensuring both prompt alignment and subject consistency across diverse styles.
Our empirical evaluations demonstrate that our method consistently outperforms prior approaches in both style and prompt alignment, while maintaining subject consistency comparable to existing methods.
@@ 3.1. Method Overview
  1. (1)
    Initial generation    We first run a vanilla generation using the SDXL model. During this pass, we store the intermediate value features $V_{sd}$, which capture the fine-grained texture and color details needed for consistent style preservation in subsequent generations.
  2. (2)
    Correspondence computation   Next, we run a generation to compute a set of DIFT features (Tang et al., 2023), which are used to establish a correspondence mapping $C_{\alpha}$ for the subject indices $s_{\alpha}$ in each image $\alpha$ and only includes attention crossing, a component which allows images to attend to subjects of other images in the self-attention layer and does not rely on the correspondence mapping. The subject indices are obtained from the cross-attention layer using a threshold over the attention map matching the subject token query. The images generated for computing DIFT do not employ the $Q$ and $K$ modifications of Sec. 3.2 since these require the subject location information that is computed using DIFT. However, the component crossing of Sec. 3.3, in which $K$ and $V$ are enriched with key and value pairs from other images in the batch after applying AdaIn is applied to obtain some level of subject consistency while maintaining style.
  3. (3)
    Final generation   In the final pass, we perform a full generation that integrates all components: i. Value-Preservation: We reuse the stored values from the initial vanilla run, maintaining stylistic consistency, ii. Attention Transfer: During the initial phase, we employ the correspondence mapping to inject query and key features, for aligning the subject details across images, and iii. Attention Crossing: Throughout the final generation, we apply attention crossing to allow image queries to attend to the Keys and Values of other images in the batch, which improves subject consistency, while using AdaIN to prevent style-leak between images due to the different Value distributions.
Figure 4. Qualitative comparison of our method with Consistory, DB-LoRA, IP-Adapter and B4M demonstrates its effectiveness across varying text descriptions, character consistency, and style alignment. Unlike other methods, our approach preserves character features and maintains consistent appearance while faithfully adhering to the specified style and textual descriptions.
@@
Table 2. Comparing our method to other zero-shot methods that achieve style or identity consistency, each with its own distinct goal. The component modification part shows the modification of the presentation of each generated image, while the component import part shows how the self-attention of the model is modified to have a cross attention component. The methods in the table are Consistory (Tewel et al., 2024), Cross-Image Attention (Alaluf et al., 2023), StyleAligned (Hertz et al., 2023) and IlluSign (Bruner et al., 2025). Our method is the only one that has an identity goal as well as a style goal, on top of the prompt faithfulness goal, which requires a much more nuanced solution. Prompt-to-Prompt (Hertz et al., 2022) is not listed as its attention modifications are done in the cross-attention layer and not the self-attention layer.
@@ 3.2. Transferring Style While Maintaining Appearance
We aim to transfer style from the target image while adopting the semantic content of the source images. To achieve this, we first perform a vanilla pass through the SDXL model, during which we store the value matrices $V_{sd}$ from the self-attention layer of the decoder at the highest-resolution transformer block $64\times 64]$, and only during the early diffusion steps, specifically at steps $\frac{n}{10}$ to $3\frac{n}{10}$ (where $n$ is the number of steps). We also obtain subject masks $s_{1},s_{2},\dots$ using a threshold over the attention maps matching the subject token query in the cross-attention layer.
@@ 3.3. AdaIN for Style Preservation in Attention Crossing
Our approach mitigates this by applying adaptive instance normalization (AdaIN) (Huang and Belongie, 2017) to the values before cross-subject attention, effectively isolating semantic content from style-specific features, since the statistical distribution of features is a key aspect of style, and texture is often defined in terms of such statistics (Haralick, 1979). By matching feature statistics using AdaIN (Huang and Belongie, 2017), we preserve the intended style: merely normalizing to constant values would distort the distribution and alter the resulting image style.
The $\mathcal{A}$ operator operation is defined as follows,
@@
where $K_{l},V_{l}$ are the keys and values in the self-attention layer matching image $l$, and $s_{l}$ are the mask indices of the subjects in the image.
@@ 3.4. Summarizing the differences from other methods
With a clear view of the method and the associated terminology, we revisit the comparison to previous work, now on a clear technical level. A comparison to the most similar contributions can be found in Table  2. The table separates the modification of the self-attention mechanism from the step of selectively importing content from other attention maps, creating a cross-attention scheme. We observe notable differences in the handling of self-attention component imports and modifications across various methods. Most approaches rely on direct Value imports from the anchor image. For instance, StyleAligned and Consistory both utilize direct Value imports, with Consistory further incorporating hidden state injections for enhanced visual consistency. Cross-Image Attention (Alaluf et al., 2023) similarly relies on direct Value modification, emphasizing precise texture and color transfer.
In contrast, our approach avoids direct Value imports to prevent the style misalignment typically associated with direct appearance transfer. Instead, we employ targeted Value modifications, aligning values to the original $SDXL$ features $V_{sd}$ for improved style alignment. Additionally, we apply AdaIN to regulate the statistical properties of imported Values, ensuring smoother integration into the target domain. Unlike Consistory and Cross-Image Attention, we do not use the hidden values of the feature embedding for the self-attention intervention, and instead modify $Q$, $K$, and $V$ selectively, giving $V$ a different treatment. StyleAligned modified only $Q$ and $K$, using AdaIn that in our method is applied for the $V$ of the dictionary expansion part (the crossing of attentions).
There are also several differences, which are not captured in Table  2. For example, the timing of the attention intervention varies across methods. Key and Query modifications as well as the $V_{sd}$ injection in Consistyle, are confined to early generation steps, while other modifications and imports are applied across all time steps as specified in Sec.  3.2. Also, the Keys and Queries modification is done only on the text-guided part of the batch in the self-attention layer, where the $V_{sd}$ injection and attention crossing components are also applied on the non-guided elements, which seems to yield a slight improvement.
@@ 4. Experiments
We evaluate our method using three core metrics: prompt alignment, consistency, and style alignment, comparing it against four state-of-the-art (SOTA) baselines designed for consistent image generation. The first baseline is a training-free Consistory approach (Tewel et al., 2024), the second is an encoder-based method, IP-Adapter (Ye et al., 2023), and the third and fourth are personalization training-based methods, DreamBooth-LoRA (DB-LoRA) (Ryu, 2023) and Break-for-Make (B4M) (Xu et al., 2025). For training details of our baseline models, please refer to the supplementary material. Unlike Consistory, which directly modifies the diffusion process, both IP-Adapter and DB-LoRA utilize a single reference image for personalization.
4.1. Implementation Details and Latency
Our method was evaluated on an A100 GPU (40GB) with a batch size of 5. SDXL and IPAdapter requires  12 sec per image. Consistory involves two passes, amounting to 24 sec per image, while our method requires three passes, totaling 36 sec per image. In contrast, training-based baselines involve substantially higher computational overhead. DB-LoRA requires around 10 minutes of training per subject. B4M entails 70 min per subject, 70 min per style, and an additional 70 min for each subject–style combination, resulting in 210 min for a single subject–style pair.
@@ 4.2. Qualitative Results
Our method is designed to accommodate a wide range of visual styles, including highly detailed photographic renderings, abstract illustrations and 3D aesthetics. In Fig.  4 and Fig.  ConsiStyle: Style Diversity in Training-Free Consistent T2I Generation we illustrate its ability to preserve style, maintain character consistency, and align with textual descriptions. As shown, B4M, IP-Adapter, and DB-LoRA exhibit strong character consistency, yet they fail to respect the intended style and often feature very little variation. Notably, in both subjects these three methods render the kitten using a realistic style and the dragon with a 3D animation style across all images, disregarding the specified stylistic variations due to overfitting to the reference image. In Fig. ConsiStyle: Style Diversity in Training-Free Consistent T2I Generation, B4M shows better alignment to the style but the other two baselines fail in this. Moreover, we observe that LoRA-based methods tend to suffer from overfitting and require an exhaustive hyperparameter search grid to maintain both content and style. Consistory demonstrates strong character consistency and faithful prompt alignment; however, it frequently fails to harmonize the subject with the stylistic setting. In Fig.  4, this issue is evident in the dragon images for the film noir and line-art styles—both intended to be black and white, yet Consistory generates colored images or subjects. This misalignment extends beyond color. For example, in the kitten images (first and fourth row), the subject exhibits a highly realistic texture that clashes with the surrounding stylistic context. Furthermore, as shown in Fig.  5, although the compositions may initially seem coherent, closer examination reveals integration flaws—characters often appear visually ‘stitched’ onto the background rather than naturally embedded within the scene. In contrast, our approach maintains both consistency and stylistic harmony, allowing characters to seamlessly integrate into their environments and preserving the intended aesthetic. The application of our method to multiple anchor images is demonstrated in Fig. 7.
@@ 4.3. Quantitative Results
Table 3. Comparison of various methods along perceptual, content, and style metrics. (Mean ± SD)
Method DreamSim $\downarrow$ CLIPScore $\uparrow$ CLIPScore, Styled $\uparrow$ LPIPS $\downarrow$ Gram L2 $\downarrow$ DINO $\uparrow$
Full Dataset (500 samples)
Consistyle (ours) $0.40\pm 0.10$ $\mathbf{32.84\pm 1.66}$ $\mathbf{36.03\pm 1.69}$ $\mathbf{0.21\pm 0.06}$ $\mathbf{1.25\pm 0.69}$ $\mathbf{0.85\pm 0.08}$
Consistory $0.33\pm 0.12$ $32.75\pm 1.53$ $35.34\pm 1.67$ $0.27\pm 0.07$ $1.85\pm 1.13$ $0.78\pm 0.11$
IP Adapter $\mathbf{0.25\pm 0.08}$ $30.98\pm 2.03$ $32.07\pm 2.08$ $0.43\pm 0.07$ $2.96\pm 1.57$ $0.54\pm 0.17$
DreamBooth-LoRA $0.28\pm 0.13$ $32.43\pm 1.74$ $34.33\pm 2.20$ $0.42\pm 0.07$ $2.60\pm 1.32$ $0.59\pm 0.15$
Cross-Img $0.55\pm 0.12$ $0.68\pm 1.36 33$$3.32\pm 1.52$ $0.33\pm 0.02$ $1.43\pm 0.70$ $0.85\pm 0.08$
IlluSign $0.53\pm 0.12$ $0.98\pm 1.48 33$$3.82\pm 1.31$ $0.38\pm 0.02$ $1.40\pm 0.67$ $0.82\pm 0.08$
Subset (100 samples)
Consistyle (ours) $0.46\pm 0.12$ $\mathbf{33.00\pm 1.60}$ $\mathbf{36.55\pm 1.18}$ $\mathbf{0.30\pm 0.04}$ $\mathbf{1.35\pm 0.9}$ $\mathbf{0.86\pm 0.08}$
B4M $\mathbf{0.38\pm 0.17}$ $0.67\pm 2.03 33$$3.30\pm 1.60$ $0.72\pm 0.02$ $2.75\pm 1.71$ $0.55\pm 0.15$
To evaluate our approach, we constructed a dataset of 25 prompt groups, each consisting of a subject description paired with ten distinct prompts. The subjects were divided into four categories: humans, animals, fantasy creatures, and inanimate objects, generated using ChatGPT. In addition, we curated two style groups, each containing ten diverse styles obtained from an online resource (Agsteiner, 2023). By combining each prompt group with each style group, we obtained 500 images, organized into 50 unique sets. Each set corresponds to one experiment and consists of a batch of $B=10$ images, where both prompts and styles vary, ensuring that no prompt or style is repeated within the same batch.
For our evaluation we employ several automated metrics following prior work (Tewel et al., 2024; Junyao et al., 2024; Avrahami et al., 2024). Text alignment is measured using CLIPScore (Hessel et al., 2021), both with and without style descriptions. Subject consistency is evaluated using DreamSim (Fu et al., 2023) as proposed by (Tewel et al., 2024) with background removal. For style alignment, we measure Gram Matrix distance (Gatys et al., 2015), and for perceptual similarity we measure LPIPS (Zhang et al., 2018), and DINO similarity (Oquab et al., 2023). These latter metrics (LPIPS and DINO) are used to evaluate content fidelity between stylized images (Xu et al., 2025; Wang et al., 2025), by comparing each stylized output to vanilla SDXL generations that serve as style references.
We conduct two sets of experiments: one using our full dataset of 500 images, and another using a subset of 100 images. The latter was necessary due to the computational cost of training B4M, which requires approximately four hours per image. The results of these experiments are presented in Table 3. As can be seen, our method outperforms all baselines in the prompt alignment and style alignment scores. However, we note a critical limitation in automated metrics for consistency when operating in style-diverse settings. Automated methods tend to over-penalize stylistic variation, potentially rewarding overly consistent outputs that lack stylistic diversity.